What is the molarity of a solution that has 0.75L of 0.25M Na2SO4


1
Expert's answer
2021-05-18T07:52:50-0400

Proportion:

0.25 mol – 1.00 L

x mol – 0.75 L

"x = \frac{0.25 \times 0.75}{1} = 0.1875 \;mol"

M(Na2SO4) = 142.04 g/mol

"m = n \times M = 0.1875 \times 142.04 = 26.63 \;g"

We need 26.63 g of Na2SO4 to prepare 0.75 L of 0.25 M Na2SO4.